An air refrigerator works between the pressure limit of 1bar and 5 bar. The temperature of the air entering the compressor and expansion cylinder is 10°C and 25°C. The expansion and compression follow the law of PV(13) C. Find the


following: 1) Theoretical COP of the refrigerating cycle


2) If the load on the refrigerating machine is 10 tons, then find the amount of air circulated per minute through the system assuming the actual COP is 50% of the theoretical COP


Assume L/D =1.5 and cp-1kJ/Kg °C and C=0.7


kJ/Kg °C for air


1
Expert's answer
2021-05-17T17:04:02-0400

1. Theoretical COP = "\dfrac{Q_C}{W}" = "\dfrac{T_C}{T_H -T_C}"


TH = 25°C = 273 + 25 = 298K

TC = 10°C = 273 + 10 = 283K


COP = "\dfrac{283}{298-283}" = "\dfrac{283}{15}" = "18.87"


2. Refrigerating effect per kg of air = cp(∆T) = 1.003(15) = 15.045 kJ/kg.


Refrigerating effect produced by the refrigerating machine:

5 × 10000 = 50000 kJ/hr.


Hence mass of air in circulation:

"\dfrac{50000}{15.045× 60}= 55.39\ kg/min"